# Improved Generalization Bounds for Communication Efficient Federated Learning

### v1

@@ Appendix C Partial Client Participation Setting
We first define an empirical risk for the partial participation distribution $\hat{\mathcal{K}}$, where $Supp(\hat{\mathcal{K}})\neq\{1,\dots,K\}$ and $|Supp(\hat{\mathcal{K}})|=\hat{K}\leq K$, on dataset $\boldsymbol{S}$ as
@@
where ${\hat{\mathcal{K}}}$ is an arbitrary distribution on participating nodes that is a part of all nodes, and $R_{\boldsymbol{S}_{k}}(M_{\boldsymbol{\theta}})$ is the empirical risk for model $M_{\boldsymbol{\theta}}$ on local dataset $\boldsymbol{S}_{k}$. We further define a population risk for model $M_{\boldsymbol{\theta}}$ as
@@
The expected generalization error is expressed as $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$. Note that the second term in (61) that is related to difference between in sample and out of sample loss can be bounded in the same way as in theorem A.2. The first term is associated with the participation of not all clients, and in the following, we demonstrate, under certain conditions, that this term would be zero in expectation.
@@
$\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}x_{k}=\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}x_{k},\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}\hat{\mathcal{K}}(k)x_{k}=\frac{K}{\hat{K}}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\mathcal{K}(k)x_{k},\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}\hat{\mathcal{K}}^{2}(k)x_{k}=\frac{K^{2}}{\hat{K}^{2}}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\mathcal{K}^{2}(k)x_{k}$ (66)
@@
So, the expected generalization error, $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$,will be just the expectation of the second term in 61 that can be bounded using Lemma A.2 by
@@
For scheme 2 we can obtain in the same way by taking the expectation of (71) with respect to $\mathcal{P}$, and considering Lemma C.2, we get the generalization bound for method $2$ as
@@ Appendix D Algorithms Used in the Experiments
In this section we have listed our implementation of FedAvg (Algorithm 2), SCAFFOLD Karimireddy et al., (2019) (Algorithm 3), and FedALS+SCAFFOLD (Algorithm 4). In Algorithm 3, we observe that in addition to the model, SCAFFOLD also keeps track of a state specific to each client, referred to as the client control variate $\boldsymbol{c}_{k,r}$. It is important to recognize that the clients within SCAFFOLD have memory and preserve the $\boldsymbol{c}_{k,r}$ and $\sum_{k=1}^{K}\boldsymbol{c}_{k,r}$ values. Additionally, when $\boldsymbol{c}_{k,r}$ consistently remains at 00, SCAFFOLD essentially becomes equivalent to FedAvg.
Algorithm 4 demonstrate the integration of FedALS and SCAFFOLD. It is important to observe that in this algorithm, the control variables are fragmented according to various model layers, and there exist distinct local step counts for different layers. This concept underlies the fundamental principle of FedALS.
@@
5:           $\boldsymbol{\theta}_{k,r,t+1}=\boldsymbol{\theta}_{k,r,t}-\eta\big{(}\frac{1}{|\mathcal{B}_{k,r,t}|}\sum_{i\in\mathcal{B}_{k,r,t}}\nabla l(M_{\boldsymbol{\theta}_{k,r,t}},\boldsymbol{z}_{{k},{i}})-\boldsymbol{c}_{k,r}+\sum_{k=1}^{K}\boldsymbol{c}_{k,r}\big{)}$
@@
7:        $\boldsymbol{c}_{k,r+1}=\sum_{k=1}^{K}\boldsymbol{c}_{k,r}-\boldsymbol{c}_{k,r}+\frac{1}{\eta\tau}(\boldsymbol{\theta}_{k,r,0}-\boldsymbol{\theta}_{k,r,\tau})$
@@
5:           $\boldsymbol{\theta}_{k,r,t+1}=\boldsymbol{\theta}_{k,r,t}-\eta\big{(}\frac{1}{|\mathcal{B}_{k,r,t}|}\sum_{i\in\mathcal{B}_{k,r,t}}\nabla l(M_{\boldsymbol{\theta}_{k,r,t}},\boldsymbol{z}_{{k},{i}})-\boldsymbol{c}_{k,r}+\sum_{k=1}^{K}\boldsymbol{c}_{k,r}\big{)}$
@@
7:              $\boldsymbol{c}_{k,r}^{\boldsymbol{h}}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}^{\boldsymbol{h}}-\boldsymbol{c}_{k,r}^{\boldsymbol{h}}+\frac{1}{\eta\tau}(\boldsymbol{h}_{k,r,0}-\boldsymbol{h}_{k,r,t})$
@@
10:              $\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}-\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}+\frac{1}{\eta\alpha\tau}(\boldsymbol{\phi}_{k,{\lfloor\frac{r\tau+t-\alpha\tau}{\tau}\rfloor},\mod(r\tau+t-\alpha\tau,\tau)}^{l}-\boldsymbol{\phi}_{k,r,t}^{l})$

### v3

@@ Appendix C Partial Client Participation Setting
We first define an empirical risk for the partial participation distribution $\hat{\mathcal{K}}$ on dataset $\boldsymbol{S}$, where $Supp(\hat{\mathcal{K}})\neq\{1,\dots,K\}$ and $|Supp(\hat{\mathcal{K}})|=\hat{K}\leq K$, as
@@
where ${\hat{\mathcal{K}}}$ is an arbitrary distribution on participating nodes that is a part of all nodes, and $R_{\boldsymbol{S}_{k}}(M_{\boldsymbol{\theta}})$ is the empirical risk for model $M_{\boldsymbol{\theta}}$ on local dataset $\boldsymbol{S}_{k}$. We further define a partial population risk for model $M_{\boldsymbol{\theta}}$ as
@@
The expected generalization error is expressed as $\operatorname{\mathbb{E}}_{{\boldsymbol{S}_{k}\sim\mathcal{D}k^{n_{k}}}{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$. Note that the second term in (62), which is related to the difference between in-sample and out-of-sample loss, can be bounded in the same way as in Theorem 4.1 and Theorem 4.3. The first term is associated with the participation of not all clients. In the following, we demonstrate that under certain conditions, this term would be zero in expectation.
@@
$\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}x_{k}=\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}x_{k},\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}\hat{\mathcal{K}}(k)x_{k}=\frac{K}{\hat{K}}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\mathcal{K}(k)x_{k},\operatorname{\mathbb{E}}_{\mathcal{P}}\operatorname{\mathbb{E}}_{k\sim{\hat{\mathcal{K}}}}\hat{\mathcal{K}}^{2}(k)x_{k}=\frac{K^{2}}{\hat{K}^{2}}\operatorname{\mathbb{E}}_{k\sim\mathcal{K}}\mathcal{K}^{2}(k)x_{k}.$ (67)
@@
The expected generalization error, $\operatorname{\mathbb{E}}_{\{\boldsymbol{S}_{k}\sim\mathcal{D}_{k}^{n_{k}}\}_{k=1}^{K}}\Delta_{\mathcal{A}}(\boldsymbol{S})$,will be just the expectation of the second term in 62 that can be bounded using Lemma A.2 by
@@
For Scheme 2, we can obtain the generalization bound in the same way by taking the expectation of (72) with respect to $\mathcal{P}$ and considering Lemma C.2. We get the generalization bound for Method 2 as:
@@ Appendix D Algorithms Used in the Experiments
In this section, we have listed our implementation of FedAvg (Algorithm 2), SCAFFOLD Karimireddy et al., (2019) (Algorithm 3), and FedALS+SCAFFOLD (Algorithm 4). In Algorithm 3, we observe that in addition to the model, SCAFFOLD also keeps track of a state-specific to each client, referred to as the client control variate $\boldsymbol{c}_{k,r}$. It is important to recognize that the clients within SCAFFOLD have memory and preserve the $\boldsymbol{c}_{k,r}$ and $\sum_{k=1}^{K}\boldsymbol{c}_{k,r}$ values. Additionally, when $\boldsymbol{c}_{k,r}$ consistently remains at 00, SCAFFOLD essentially becomes equivalent to FedAvg.
Algorithm 4 demonstrates the integration of FedALS and SCAFFOLD. It is important to observe that in this algorithm, the control variables are fragmented according to different model partitions as distinct local step counts exist for different parts.
@@
5:           $\boldsymbol{\theta}_{k,r,t+1}=\boldsymbol{\theta}_{k,r,t}-\eta\big{(}\frac{1}{|\mathcal{B}_{k,r,t}|}\sum_{i\in\mathcal{B}_{k,r,t}}\nabla l(M_{\boldsymbol{\theta}_{k,r,t}},\boldsymbol{z}_{{k},{i}})-\boldsymbol{c}_{k,r}+\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}\big{)}$
@@
7:        $\boldsymbol{c}_{k,r+1}=\boldsymbol{c}_{k,r}-\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}+\frac{1}{\eta\tau}(\boldsymbol{\theta}_{k,r,0}-\boldsymbol{\theta}_{k,r,\tau})$
@@
5:           $\boldsymbol{\theta}_{k,r,t+1}=\boldsymbol{\theta}_{k,r,t}-\eta\big{(}\frac{1}{|\mathcal{B}_{k,r,t}|}\sum_{i\in\mathcal{B}_{k,r,t}}\nabla l(M_{\boldsymbol{\theta}_{k,r,t}},\boldsymbol{z}_{{k},{i}})-\boldsymbol{c}_{k,r}+\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}\big{)}$
@@
7:              $\boldsymbol{c}_{k,r}^{\boldsymbol{h}}\leftarrow\boldsymbol{c}_{k,r}^{\boldsymbol{h}}-\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}^{\boldsymbol{h}}+\frac{1}{\eta\tau}(\boldsymbol{h}_{k,r,0}-\boldsymbol{h}_{k,r,t})$
@@
10:              $\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}\leftarrow\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}-\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{c}_{k,r}^{\boldsymbol{\phi}}+\frac{1}{\eta\alpha\tau}(\boldsymbol{\phi}_{k,{\lfloor\frac{r\tau+t-\alpha\tau}{\tau}\rfloor},\mod(r\tau+t-\alpha\tau,\tau)}^{l}-\boldsymbol{\phi}_{k,r,t}^{l})$